ADMISSIBILITY OF PREDICTION INTERVALS

YOSHIKAZU TAKADA

Department of Mathematics, Faculty of Science, Kumamoto University, Kumamoto 860, Japan

(Received July 15, 1993; revised June 10, 1994)

Abstract. Admissibility of prediction intervals is considered in a specified family. It is shown that the best invariant prediction interval is strongly admissible in a location family and in a scale family. Though the similar result has not been obtained for a location and scale family, the best invariant prediction interval for a normal distribution is shown to be weakly admissible.

Key words and phrases: Prediction interval, admissibility, invariance, location family, scale family, location and scale family.